OpenStudy (anonymous):

The perimeter of a rectangle is 96 ft. Find the dimensions of the rectangle if the ratio of the length to the width is 7: 5. Which of the following would be the best equation to use to solve this problem? (Points : 1)

OpenStudy (anonymous):

2x+2x=96 7x+5X=96 2(7x)+2(5X)=96 7x/5x=96 < fraction btw